A water treatment plant is designed to remove 12.0 μm diameter particles with a density of 1.2 g/mL. The water is at 10°C. The plant takes in water at a rate of 0.10 m3/s. The rectangular sedimentation tank has a depth of 4 m and a length:width ratio of 5:1. Calculate the basin’s minimum width (in m) required to remove the particles.(Show solution)


1
Expert's answer
2021-05-10T01:51:15-0400

Volume of tank= 4 × 3.33 × 0.66 = 8.79m3

Minimum width = "\frac{1 }{6}×4=\frac{2}{3}m"